# A Survey on 3D Gaussian Splatting

### v4

@@ 3.2 Optimization of 3D Gaussian Splatting
At the heart of 3D GS lies an optimization procedure devised to construct a copious collection of 3D Gaussians that accurately captures the scene’s essence, thereby facilitating free-viewpoint rendering. On the one hand, the properties of 3D Gaussians should be optimized via differentiable rendering to fit the textures of a given scene. On the other hand, the number of 3D Gaussians that can represent a given scene well is unknown in advance. One promising avenue is to let the neural network automatically learn the density of 3D Gaussians. We will introduce how to optimize the properties of each Gaussian in Sec. 3.2.1 and how to control the density of the Gaussians in Sec. 3.2.2. The two procedures are interleaved within the optimization workflow. Since there are many manually set hyperparameters in the optimization process, we omit the notations of most hyperparameters for clarity.
@@ 3.2.1 Parameter Optimization
where $\lambda\in[0,1]$ is a weighting factor. The loss function of 3D GS is slightly different from that of NeRFs. NeRFs typically calculate the loss at the pixel level rather than the image level because of the costly ray-marching.
$\bullet$ Parameter Update. Most properties of a 3D Gaussian can be optimized directly through back-propagation. It is essential to note that directly optimizing the covariance matrix $\bm{\Sigma}$ can result in a non-positive semi-definite matrix, which would not adhere to the physical interpretation typically associated with covariance matrices. To circumvent this issue, 3D GS chooses to optimize a quaternion $\bm{q}$ and a 3D vector $\bm{s}$. $\bm{q}$ and $\bm{s}$ represent rotation and scale, respectively. This approach allows the covariance matrix $\bm{\Sigma}$ to be reconstructed as follows:
@@
where $\bm{R}$ and $\bm{S}$ denote the rotation and scaling matrix derived from $\bm{q}$ and $\bm{s}$, respectively. There is a complex computational graph to obtain the opacity $\alpha$, i.e., $\bm{q}$ and $\bm{s}\mapsto\bm{\Sigma}$, $\bm{\Sigma}\mapsto\bm{\Sigma}^{\prime}$, and $\bm{\Sigma}^{\prime}\mapsto\alpha$. To avoid the cost of automatic differentiation, 3D GS derives the gradients for $\bm{q}$ and $\bm{s}$ so as to compute them directly during optimization.
@@ 3.2.2 Density Control
Though 3D GS has achieved impressive milestones, significant room for improvement remains, e.g., data and hardware requirement, rendering and optimization algorithm, and applications in downstream tasks. In the subsequent sections, we seek to elaborate on select extended versions of 3D GS. These are: i) Data-efficient 3D GS [45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55] (Sec. 4.1), ii) Memory-efficient 3D GS [56, 57, 58, 59, 60, 61, 62, 63, 64] (Sec. 4.2), iii) Photorealistic 3D GS [65, 66, 67, 68, 69, 70, 71, 72, 73, 74, 75, 76, 77, 78, 79, 80] (Sec. 4.3), iv) Improved Optimization Algorithms [81, 22, 82, 77, 83, 84, 85, 86] (Sec. 4.4), v) 3D Gaussian with More Properties [87, 88, 89, 90, 91, 92, 93] (Sec. 4.5), and vi) 3D GS with Structured Information [94, 95, 96] (Sec. 4.6).
4.1 Data-efficient 3D GS
@@
Two primary strategies are used for data-efficient 3D GS. i) Regularization based methods introduce additional constraints such as depth information to enhance the detail and global consistency [49, 46, 51, 55]. For example, DNGaussian [49] introduced a depth-regularized approach to address the challenge of geometry degradation in sparse input views. FSGS [46] devised a Gaussian Unpooling process for initialization and also introduced depth regularization. MVSplat [51] proposed a cost volume representation so as to provide geometry cues. Unfortunately, when dealing with a limited number of views, or even just one, the efficacy of regularization techniques tends to diminish, which leads to ii) generalizability based methods that focus specially on learning priors [47, 48, 53]. One typical implementation is to generate 3D Gaussians that can be used for rendering directly without optimization, by using deep neural networks. This paradigm typically requires multiple views for training but can reconstruct the 3D scene with only one input image. For instance, PixelSplat [47] proposed to sample Gaussians from dense probability distributions. It incorporated a multi-view epipolar transformer and a reparameterization trick to avoid local minima and maintain gradient flow. Splatter Image [48] applied GS in a monocular setting through a learning-based approach, utilizing a 2D image-to-image network that maps an input image to a 3D Gaussian per pixel. Note that this paradigm is mainly focused on the reconstruction of objects, and its generalizability leaves a lot of room for improvement.
@@ 4.2 Memory-efficient 3D GS
There are two main directions to reduce memory usage. i) The first involves reducing the number of 3D Gaussians [63, 58, 62], i.e., pruning insignificant 3D Gaussians. For example, Papantonakis et al. [63] proposed a resolution-aware pruning method to prune Gaussians, reducing the Gaussian count by half. Lee et al. [58] introduced a new volume-based masking strategy that efficiently reduces the quantity of Gaussians without compromising performance. ii) The second category focuses on compressing the memory usage of 3D Gaussian properties [61, 58, 62]. For instance, Niedermayr et al. [61] compressed color and Gaussian parameters into compact codebooks, using sensitivity measures for effective quantization and fine-tuning. HAC [62] predicted the probability of each quantized attribute using Gaussian distributions and then devise an adaptive quantization module. Although current methods achieve compression ratios of several to dozens of times for storage (after training), there remains considerable potential for reducing memory usage during the training phase.
@@ 4.3 Photorealistic 3D GS
The current rendering pipeline of 3D GS (Sec. 3.1) is straightforward and involves several drawbacks. For instance, the simple visibility algorithm may lead to a drastic switch in the depth/blending order of Gaussians [10]. The realness of rendered images, including aspects such as aliasing, reflections, and artifacts, can be further optimized.
Here we list several key points for enhancing realness. i) Varying Resolutions [67, 78]. Due to the discrete sampling paradigm (viewing each pixel as a single point instead of an area), 3D GS is susceptible to aliasing when dealing with varying resolutions, which leads to blurring or jagged edges. Yan et al. [67] argued that this is mainly because traditional rendering methods cannot effectively manage the disparity between the sampling frequency of pixels and the high-frequency details of the scene, leading to visual artifacts and performance issues. Hence, they introduced multi-scale 3D GS, where the scene is represented using Gaussians of varying sizes. Analytic-Splatting [78] adopted an analytic approximation of the Gaussian integral within the pixel area, leveraging a conditioned logistic function for the cumulative distribution function to better capture the intensity response of pixels. ii) Reflections [68, 97, 98]. Achieving realistic rendering of reflective materials is a hard and long-standing problem in 3D scene reconstruction. GaussianShader [68] enhanced neural rendering for scenes with reflective surfaces by integrating a simplified shading function with 3D Gaussians. iii) Geometry. One limitation of 3D GS is the neglect of underlying scene geometry and structure, particularly in complex scenes and under varying view and lighting conditions. This sparks research on geometry-aware reconstruction [99, 100, 44, 22, 101, 102]. For example, GeoGaussian [77] focused on preserving the geometry of non-textured regions like walls and furniture, which tend to degrade over time.
@@ 4.4 Improved Optimization Algorithms
The anisotropic Gaussians, while beneficial for representing complex geometries, can create undesirable visual artifacts. For example, those large 3D Gaussians, especially in regions with view-dependent appearance, can cause popping artifacts, where visual elements abruptly appear or disappear, breaking the immersion. Further, incorporating additional regularization (e.g., geometry [77, 83] and frequency [84]) and improving the optimization process of 3D GS (Sec. 3.2) may accelerate convergence, smooth visual noise, and improve quality of rendered images.
Three main directions stand out for improving the optimization of 3D GS. i) Introducing Additional Regularization [84, 22]. 3D GS often faces challenges with over-reconstruction, where sparse, large 3D Gaussians cause blur and artifacts due to poor representation in high-variance regions. To address this, FreGS [84] introduced a progressive frequency regularization method, which refines Gaussian densification from a frequency perspective. Another notable branch is geometry-aware reconstruction, as introduced in Sec. 4.3. This line of work focuses in particular on how to preserve scene’s structure. For instance, Scaffold-GS [22] introduced a sparse grid of anchor points to organize local 3D Gaussians, which are dynamically adjusted for attributes like opacity and color based on the viewer’s perspective and distance. ii) Improving Optimization Procedure [44, 77]. In addressing the challenges of dense initialization for texture-less surfaces, especially in large-scale scenes, GaussianPro [44] devised an advanced Gaussian densification (§3.2.2) strategy using the priors of reconstructed geometries and patch matching techniques. iii) Relaxing Constraints in Optimization [81, 82]. Reliance on external tools/algorithms for can introduce errors and cap the system’s performance potential. For instance, SfM, commonly used in the initialization process, is error-prone and struggle with complex scenes. Yang et al. [81] proposed COLMAP-Free 3D GS, which introduces video stream continuity and an explicit point cloud representation so as to eliminate the need for SfM preprocessing. Though impressive, existing methods primarily concentrate on optimizing Gaussians to accurately reconstruct scenes from scratch, neglecting a challenging yet promising paradigm which reconstructs scenes in a few-shot manner through established “meta representations”. See “learning physical priors from large-scale data” in Sec. 7 for further insights.
@@
[104], [105], [106], [107], [108], [109], [110],[111], [112], [113], [114], [115], [116], [117][118], [94], [95], [93], [119], [120],[121], [122], [103], [123], [124], [125][126], [127], [128], [129], [130], [131],[132], [133], [134], [135], [90], [136][137], [138], [139][140], [141], [142][143], [116], [144][44], [145], [146], [147], [148], [149], [150], [151], [152]
Figure 6: Typical applications that have been benefited from GS (Sec. 5). Some images are borrowed from [125, 128, 138, 142, 151] and redrawn.
4.6 3D GS with Structured Information
@@ 5 Application Areas and Tasks
Building on the rapid advancements in 3D GS, a wide range of innovative applications has emerged across multiple domains (Fig. 6) such as robotics (Sec. 5.1), dynamic scene reconstruction and representation (Sec. 5.2), AI-generated content (Sec. 5.3), autonomous driving (Sec. 5.4), medical systems (Sec. 5.5), large-scale scene reconstruction (Sec. 5.6), and even other scientific disciplines [154, 155, 24, 156]. Here, we highlight key examples that underscore the transformative impact and potential of 3D GS.
5.1 Simultaneous Localization and Mapping (SLAM)
SLAM is a computational problem central to robotics and autonomous systems. It involves the challenge of a robot/device understanding its position in an unknown environment while simultaneously mapping the environment’s layout. SLAM is critical in various applications such as autonomous vehicles and robotic navigation. The core of SLAM is to create a map of an unknown environment and determine the device’s location on this map in real-time. As a result, SLAM poses great challenges for computationally intensive scene representation techniques, yet serves as a good testbed for 3D GS [157].
3D GS enters the SLAM domain as an innovative approach for scene representation. Traditional SLAM systems often use point/surfel clouds or voxel grids to represent environments [158, 159, 160, 161, 162, 163, 164, 165]. In contrast, 3D GS utilizes anisotropic Gaussians to better represent environments. Recent innovative studies have employed 3D GS in SLAM [108, 106, 107, 104, 105, 166, 110, 113], showcasing the potential and versatility of this paradigm. An intuitive way is to use 3D Gaussian as the basic representation of the dense map and optimize the tracking process. For instance, GS-SLAM [104] adopted an adaptive strategy to add or remove 3D Gaussians so as to optimize scene geometry reconstruction and improve the mapping of previously observed areas. Simiarly, Sun et al. [113] proposed additional regularization terms during the online mapping to avoid overfiting to the latest frame. SplaTAM [105] integrated a straightforward online tracking and mapping method that leverages a silhouette mask for capturing the presence of scene density, facilitating dense optimization and structured map expansion. On the other hand, designing advanced scene representations is worth exploring. For example, Photo-SLAM [108] proposed a hybrid Gaussian map, which combines explicit geometric features for precise localization with implicit photometric features for texture mapping. In addition, one interesting problem is how the Gaussian based map representations can help robotics tasks, and some early attempts in navigation have been made [167, 168, 169]. Though impressive, existing GS based SLAM systems still face persistent challenges such as dynamic elements, sensor noise, non-Lambertian objects, and depth ambiguities that require further exploration.

### v5

@@ 3.2 Optimization of 3D Gaussian Splatting
At the heart of 3D GS lies an optimization procedure devised to construct a copious collection of 3D Gaussians that accurately captures the scene’s essence, thereby facilitating free-viewpoint rendering. On the one hand, the properties of 3D Gaussians should be optimized via differentiable rasterization to fit the textures of a given scene. On the other hand, the number of 3D Gaussians that can represent a given scene well is unknown in advance. We will introduce how to optimize the properties of each Gaussian in Sec. 3.2.1 and how to adaptively control the density of the Gaussians in Sec. 3.2.2. The two procedures are interleaved within the optimization workflow. Since there are many manually set hyperparameters in the optimization process, we omit the notations of most hyperparameters for clarity.
5
@@ 3.2.1 Parameter Optimization
where $\lambda\in[0,1]$ is a weighting factor.
$\bullet$ Parameter Update. Most properties of a 3D Gaussian can be optimized directly through back-propagation. It is essential to note that directly optimizing the covariance matrix $\bm{\Sigma}$ can result in a non-positive semi-definite matrix, which would not adhere to the physical interpretation typically associated with covariance matrices. To circumvent this issue, 3D GS chooses to optimize a quaternion $\bm{q}$ and a 3D vector $\bm{s}$. Here $\bm{q}$ and $\bm{s}$ represent rotation and scale, respectively. This approach allows the covariance matrix $\bm{\Sigma}$ to be reconstructed as follows:
@@
where $\bm{R}$ is the rotation matrix derived from the quaternion $\bm{q}$, and $\bm{S}$ is the scaling matrix given by $\text{diag}(\bm{s})$. As seen, there is a complex computational graph to obtain the opacity $\alpha$, i.e., $\bm{q}$ and $\bm{s}\mapsto\bm{\Sigma}$, $\bm{\Sigma}\mapsto\bm{\Sigma}^{\prime}$, and $\bm{\Sigma}^{\prime}\mapsto\alpha$. To avoid the cost of automatic differentiation, 3D GS derives the gradients for $\bm{q}$ and $\bm{s}$ so as to compute them directly during optimization.
@@ 3.2.2 Density Control
Though 3D GS has achieved impressive milestones, significant room for improvement remains, e.g., data and hardware requirement, rendering and optimization algorithm, and applications in downstream tasks. In the subsequent sections, we seek to elaborate on select extended versions. These are: i) 3D GS for Sparse Input [45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55] (Sec. 4.1), ii) Memory-efficient 3D GS [56, 57, 58, 59, 60, 61, 62, 63, 64] (Sec. 4.2), iii) Photorealistic 3D GS [65, 66, 67, 68, 69, 70, 71, 72, 73, 74, 75, 76, 77, 78, 79, 80] (Sec. 4.3), iv) Improved Optimization Algorithms [81, 22, 82, 77, 83, 84, 85, 86] (Sec. 4.4), v) 3D Gaussian with More Properties [87, 88, 89, 90, 91, 92, 93] (Sec. 4.5), vi) Hybrid Representation [94, 95, 96] (Sec. 4.6), and vii) New Rendering Algorithm (Sec. 4.7). While we have carefully selected several key directions, we acknowledge that it is inevitably a biased view. A more comprehensive collection is given in Github.
4.1 3D GS for Sparse Input
@@
Existing methods can be categorized into two primary groups. i) Regularization based methods introduce additional constraints such as depth information to enhance the detail and global consistency [49, 46, 51, 55]. For example, DNGaussian [49] introduced a depth-regularized approach to address the challenge of geometry degradation in sparse input. FSGS [46] devised a Gaussian Unpooling process for initialization and also introduced depth regularization. MVSplat [51] proposed a cost volume representation so as to provide geometry cues. Unfortunately, when dealing with a limited number of views, or even just one, the efficacy of regularization techniques tends to diminish, which leads to ii) generalizability based methods that use learned priors [47, 48, 53, 97]. One approach involves synthesizing additional views through generative models, which can be seamlessly integrated into existing reconstruction pipelines [98]. However, this augmentation strategy is computationally intensive and inherently bounded by the capabilities of the used generative model. Another well-known paradigm employs feed-forward Gaussian model to directly generates the properties of a set of 3D Gaussians. This paradigm typically requires multiple views for training but can reconstruct 3D scenes with only one input image. For instance, PixelSplat [47] proposed to sample Gaussians from dense probability distributions. Splatter Image [48] introduced a 2D image-to-image network that maps an input image to a 3D Gaussian per pixel. However, as the generated pixel-aligned Gaussians are distributed nearly evenly in the space, they struggle to represent high-frequency details and smoother regions with an appropriate number of Gaussians.
The challenge of 3D GS for sparse inputs centers on the modeling of priors, whether through depth information, generative models, or feed-forward Gaussian models. The fundamental trade-off lies between overfitting to available views and using learned priors for generalization. Future research could explore adaptive mechanisms for controlling this trade-off, potentially through learned confidence measures, context-aware prior selection, user preferences, etc. In addition, while current methods focus on static scenes, extending these approaches to dynamic scenarios presents an exciting frontier for investigation, particularly in handling temporal consistency and motion-induced artifacts.
@@ 4.2 Memory-efficient 3D GS
Recent research has pursued two primary directions to address memory efficiency. First, several approaches focus on reducing the number of 3D Gaussians [63, 58, 62]. These methods either employ strategic pruning of low-impact Gaussians, such as the volume-based masking [58], or represent neighboring Gaussians using the same properties stored within a “local anchor” obtained by clustering [22], hash-grid [62], etc. Second, researchers have developed methods for compressing Gaussian’s properties [61, 58, 62]. For instance, Niedermayr et al. [61] compressed color and Gaussian parameters into compact codebooks, using sensitivity measures for effective quantization and fine-tuning. HAC [62] predicted the probability of each quantized attribute using Gaussian distributions and then devise an adaptive quantization module. These directions are not mutually exclusive; instead, one framework might use a hybrid approach combining multiple strategies.
While current compression techniques have achieved significant storage reduction ratios (often by factors of 10-20$\times$), several challenges remain. The field particularly needs advances in memory efficiency during the training phase, potentially through quantization-aware training protocols, the development of scene-agnostic, reusable codebooks, etc. Furthermore, optimizing the trade-off between compression efficiency and visual fidelity remains an open problem.
@@ 4.3 Photorealistic 3D GS
The current rendering pipeline of 3D GS (Sec. 3.1) is straightforward and involves several drawbacks. For instance, the simple visibility algorithm may lead to a drastic switch in the depth/blending order of Gaussians [10]. The visual fidelity of rendered images, including aspects such as aliasing, reflections, and artifacts, can be further optimized.
Recent research has focused on addressing three main aspects of visual quality, with aliasing being specific to 3D GS’s rendering algorithm, while reflection and blur handling represent broader challenges in 3D reconstruction. i) Aliasing. Due to the discrete sampling paradigm (viewing each pixel as a single point instead of an area), 3D GS is susceptible to aliasing when dealing with varying resolutions, which leads to blurring or jagged edges. Solutions emerged at both training and inference stages. Researchers developed training-time improvements from the sampling rate perspective and introduced schemes such as multi-scale Gaussians [67], 2D Mip filter [65], and conditioned logistic function [78]. Inference-time solutions, such as 2D scale-adaptive filtering [80], offer enhanced fidelity that can be integrated into any existing 3D GS frameworks. ii) Reflection. Achieving realistic rendering of reflective materials is a hard, long-standing problem in 3D scene reconstruction. Recent works have introduced various approaches to model reflective materials [68, 99, 73] and enable relightable Gaussian representation [23], though achieving physically accurate specular effects remains challenging. iii) Blur. While 3D GS excels on carefully curated datasets, real-world captures often suffer from blurs such as motion blur and defocus blur. Recent approaches explicitly incorporated blur modeling during training, employing techniques such as coarse-to-fine kernel optimization [74] and photometric bundle adjustment [75] to address this challenge.
While the approximations made in 3D GS (Sec. 3.1) contribute to its computational efficiency, they also lead to aliasing, difficulties in illumination estimation, etc. Current solutions, though impressive, typically address individual problems rather than providing a universal solution. A practical intermediate approach involves first detecting specific issues (e.g., aliasing, blur) and then applying targeted optimization strategies. The ultimate goal remains developing an advanced reconstruction system that overcomes these limitations, either through fundamental improvements to 3D GS or through brand-new architectures.
@@ 4.4 Improved Optimization Algorithms
The optimization of 3D GS presents several challenges that affect the quality of reconstruction. These include issues with convergence speed, visual artifacts from improper Gaussians, and the need for better regularization during optimization. The raw optimization method (Sec. 3.2) might lead to overreconstruction in some regions while underrepresenting others, resulting in blur and visual inconsistencies.
Three main directions stand out for improving the optimization of 3D GS. i) Additional Regularization (e.g., frequency [84] and geometry [22, 77]). Geometry-aware approaches have been particularly successful, preserving scene structure through the incorporation of local anchor points [22], depth and surface constraints [100, 101, 102], Gaussian volumes [103], etc. ii) Optimization Procedure Enhancement [44, 101, 104]. While the original strategy of density control (Sec. 3.2.2) has proven valuable, considerable room for improvement remains. For example, GaussianPro [44] addresses the challenge of dense initialization in texture-less surfaces and large-scale scenes through an advanced Gaussian densification strategy. iii) Constraint Relaxation. Reliance on external tools/algorithms can introduce errors and cap the system’s performance potential. For instance, SfM, commonly used in the initialization process, is error-prone and struggle with complex scenes. Recent works have begun exploring COLMAP-free approaches utilizing stream continuity [81, 105], potentially enabling learning from internet-scale unposed video datasets.
Though impressive, existing methods primarily concentrate on optimizing Gaussians to accurately reconstruct scenes from scratch, neglecting a challenging yet promising solution which reconstructs scenes in a few-shot manner through established “meta representations”. Such solution could enable adaptive meta-learning strategies that combine scene-specific and general knowledge. See “learning physical priors from large-scale data” in Sec. 7 for further insights.
@@
4.6 Hybrid Representation
@@
4.7 New Rendering Algorithm for 3D Gaussians
While the rasterization-based pipeline of 3D GS offers impressive real-time performance, it still suffers from the inherent limitations, including inefficient handling of highly-distorted cameras (crucial for robotics), secondary rays (for optical effects like reflections and shadows), and stochastic ray sampling (needed in various existing pipelines). In addition, the assumptions that Gaussians do not overlap and can be sorted accurately using only centers are often violated in practice, leading to temporal artifacts when camera movement changes sorting order.
Recent works [108, 109, 110] explored ray tracing based rendering algorithms as an alternative. For instance, GaussianTracer [108] introduced a new ray tracing implementation for Gaussian primitives, and devised several accelerating strategies according to the uneven density and interleaved nature of Gaussians. EVER [109] deivsed a physically accurate, constant density ellipsoid representation that allows for the exact computation of the volume rendering integral, rather than relying on somewhat satisfactory approximations. This advancement eliminates popping artifacts.
Thanks to the fundamental paradigm shift, several exciting possibilities might emerge, including advanced optical effects (reflection, refraction, shadows, global illumination, etc.), support for complex camera models (highly-distorted lenses, rolling shutter effects, etc.), physically accurate rendering with true directional appearance evaluation (vs. tile based approximation), and more. While these capabilities currently come with additional computational costs, they provide essential building blocks for future research in inverse rendering, physical material modeling, relighting, and complex scene reconstruction.
5.15.25.35.45.55.65.7[111], [112], [113], [114], [115], [116], [117],[118], [119], [120], [121], [122], [123], [124][125], [94], [95], [93], [126], [127],[128], [129], [106], [130], [131], [132][133], [134], [135], [136], [137], [138],[139], [140], [141], [142], [90], [143][144], [145], [146], [147][148], [149], [150], [151][152], [153], [154][155], [156], [157][158], [159], [160], [161][44], [162], [163], [164], [165][166], [167], [168], [143], [169][170], [21], [171], [172], [173]
Figure 6: Typical applications benefited from GS (Sec. 5). Some images are borrowed from [132, 135, 154, 160, 146, 166] and redrawn.
@@ 5 Application Areas and Tasks
Building on the rapid advancements in 3D GS, a wide range of innovative applications has emerged across multiple domains (Fig. 6) such as robotics (Sec. 5.1), dynamic scene reconstruction and representation (Sec. 5.2), generation and editing (Sec. 5.3), avatar (Sec. 5.4), medical systems (Sec. 5.5), large-scale scene reconstruction (Sec. 5.6), physics (Sec. 5.7), and even other scientific disciplines [174, 175, 24, 176]. Here, we highlight key examples that underscore the transformative impact and potential of 3D GS and offer a more comprehensive collection in Github.
5.1 Robotics
The evolution of scene representation in robotics has been profoundly shaped by the emergence of NeRF, which revolutionized dense mapping and environmental interaction through implicit neural models. However, NeRF’s computational cost poses a critical bottleneck for real-time robotic applications. The shift from implicit to explicit representation not only accelerates optimization but also unlocks direct access to spatial and structural scene data, making 3D GS a transformative tool for robotics. Its ability to balance high-fidelity reconstruction with computational efficiency positions 3D GS as a cornerstone for advancing robotic perception, manipulation, and navigation in dynamic, real-world environments.
The integration of GS into robotic systems has yielded significant advancements across three core domains. In SLAM, GS-based methods [115, 113, 114, 111, 112, 177, 117, 178, 179, 116, 123, 180, 181, 182, 124] excel in real-time dense mapping but face inherent trade-offs. Visual SLAM frameworks, particularly RGB-D variants [112, 114, 178], leverage depth supervision for geometric fidelity but falter in low-texture or motion-degraded environments. RGB-only approaches [113, 115, 183] circumvent depth sensors but grapple with scale ambiguity and drift. Multi-sensor fusion strategies, such as LiDAR integration [177, 182, 159], enhance robustness in unstructured settings at the cost of calibration complexity. Semantic SLAM [123, 117, 116] extends scene understanding through object-level semantics but struggles with scalability due to lighting sensitivity in color-based methods or computational overhead in feature-based methods. 3D GS based manipulation [118, 119, 120, 121, 122] bypasses the need for auxiliary pose estimation in NeRF-based methods, enabling rapid single-stage tasks like grasping in static environments via geometric and semantic attributes encoded in Gaussian properties. Multi-stage manipulation [118, 120], where environmental dynamics demand real-time map updates, requires explicit modeling of dynamic adjustments (e.g., object motions and interactions), material compliance, etc.
The advancement of 3D GS in robotics faces three pivotal challenges. First, adaptability in dynamic and unstructured environments remains critical: real-world scenes are rarely static, requiring systems to continuously update representations amid motion, occlusions, and sensor noise without sacrificing accuracy. Second, current semantic mapping methods rely on costly, scene-specific optimization processes, limiting generalizability and scalability for real-world deployment. Third, unlike NeRF based systems which can use MLP parameters as input features for downstream decision-making, 3D Gaussians’ inherent lack of spatial order complicates feature aggregation, with no standardized framework yet established. Bridging the gap between high-fidelity reconstruction and actionable semantic/physical understanding will define the next frontier for 3D GS, moving beyond passive mapping towards embodied intelligence.